Paymaster Enterprises has arranged to finance its seasonal? working-capital needs with a? short-term bank loan. The loan will carry a rate of 14 percent per annum with interest paid in advance? (discounted). In? addition, Paymaster must maintain a minimum demand deposit with the bank of 11 percent of the loan balance throughout the term of the loan. If Paymaster plans to borrow $90,000 for a period of 4 ?months, what is the annualized cost of the bank? loan? (Round to 2 decimal places)
Answer #1

Interest expense = 0.14*90000*4/12 =$4200

Assuming that Paymaster has to leave 11% of the loan idle in a compensating balance the annualized cost of the bank loan can be calculated as follow:

11% of Loan =0.11*90000 = $9900, Interest amount will deduct from loan (Discount loan)

annualized cost of the bank loan = 4200*12/(90000-9900-4200)*4 = 12600/75900

Annualized cost of the bank loan = 16.60%